OpenStudy (anonymous):

One pipe fills a tank in 1 hour while another pipe can fill the same tank in 2 hours. How long will it take to fill the tank if both pipes are used

hero (hero):

\(\dfrac{1\times 2}{2 + 1} = t\) \(t = \text{hours}\)

hero (hero):

\(\dfrac{2}{3} = t\) Which basically means that the job will be done in 2/3 of an hour

hero (hero):

You can convert hours to minutes in this case for it to make more sense. \(\dfrac{2}{3}(60 \space\text{min}) = 40 \space\text{min}\)